OpenStudy (anonymous):

HELP PLEASE!!!! find the equation of ellipse. center @origin. Passing through (2, square root 30 / 3.), distance between directrices 24 square root 7 / 7.

OpenStudy (anonymous):

\[\frac{ x^{2} }{ a^{2} }+\frac{ y ^{2} }{ b^{2}}=1\] can you identify this equation?

OpenStudy (anonymous):

yes its ellipse

OpenStudy (anonymous):

so plug in the values of (x,y) i.e \[(2,\sqrt{30}/3)\] in that equation

OpenStudy (anonymous):

okayy..

OpenStudy (anonymous):

now dist. of directrix from origin is a/e where e^2=1-(b/a)^2

OpenStudy (anonymous):

so dist between 2 directrices will be 2a/e

OpenStudy (anonymous):

yes and eccentricity not given...

OpenStudy (anonymous):

so we find eccentricity in terms of a^2 and b^2

OpenStudy (anonymous):

how do we get a? b?

OpenStudy (anonymous):

okay let's visualize the problem first

OpenStudy (anonymous):

you want the equation right?

OpenStudy (anonymous):

so you want a, b

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

so first lets use the co-ordinates given and plug it in the master eqn of ellipse

OpenStudy (anonymous):

so you get a eqn with 2 unknown i.e a,b

OpenStudy (anonymous):

you need 1 more to find the exact values for a,b

OpenStudy (anonymous):

for that 2a/e = dist between directrices

OpenStudy (anonymous):

2 coordinates?

OpenStudy (anonymous):

itz co-ordinate not plural ..sorry

OpenStudy (anonymous):

typo..

OpenStudy (anonymous):

co-ordinate is (2, sqrt30/3) use it in ellipse eqn

OpenStudy (anonymous):

eqn will be 4/a^2 +10/b^2 =1

OpenStudy (anonymous):

yeah. thats what i did

OpenStudy (anonymous):

now you answer me what's 2a/e?

OpenStudy (anonymous):

distance between directrix and i dont know how to get e

OpenStudy (anonymous):



OpenStudy (anonymous):

\[e ^{2}=1-\ \frac{ b ^{2} }{ a ^{2}}\]

OpenStudy (anonymous):

therefore plug in this eqn to find dist .. i.e \[4a ^{2}/e ^{2}=dist ^{2}\]

OpenStudy (anonymous):

where dist =

OpenStudy (anonymous):

24 sqrt 7/7

OpenStudy (anonymous):

4a^2 / e^2.. how did you come up with this??

OpenStudy (anonymous):

look at the graph directix lies at a/e.. therefore dist between both of them is 2a/e

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yey!

OpenStudy (anonymous):

yippie!!

OpenStudy (anonymous):

and.. how about getiing e?

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

i already ut the dist on the equation and what's next?

OpenStudy (anonymous):

already put the distance on the equation

OpenStudy (anonymous):

e defined like that.. and for ellipse its the eqn written above ingeneral e=distance of locus of point from focus/ perpendicular dist. of locus point from directrix

OpenStudy (anonymous):

now you get a eqn in a^2 and b^2 again

OpenStudy (anonymous):

right??

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

so you can now solve for 2 linear eqn with 2 variable .. right..?? note that you need to find a^2 and b^2 not a, b.. so i used the term linear eqn as you can put a^2=p and b^2=q

OpenStudy (anonymous):

still confused?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

at which part?

OpenStudy (anonymous):

the e part.. dont know how to solve. what i know is the first equation would be the 4/a^2 +10/b^2 =1 and yeah stuck up no idea about the secodn equation..

OpenStudy (anonymous):

suppose u have a eqn ... 2a/e=24 sqrt7/7

OpenStudy (anonymous):

now u have a eqn of e in terms of a^2,b^2.. (the eqn just below the graph i drew)

OpenStudy (anonymous):

now substitute this eqn of e in the dist eqn above..

OpenStudy (anonymous):

square both sides.. u'll get eqn. in a^2 and b^2

OpenStudy (anonymous):

and..

OpenStudy (anonymous):

so now substitute a^2 by p and b^2 by q in both the obtained .. u get eqn in p,q .. now from one eqn find p in terms of q .. then substitute this in the second equation